\begin{thebibliography}{31}
\expandafter\ifx\csname natexlab\endcsname\relax\def\natexlab#1{#1}\fi

\bibitem[{Beltagy et~al.(2019)Beltagy, Lo, and Cohan}]{scibert}
Iz~Beltagy, Kyle Lo, and Arman Cohan. 2019.
\newblock \href {https://doi.org/10.18653/v1/D19-1371} {{S}ci{BERT}: A
  pretrained language model for scientific text}.
\newblock In \emph{Proceedings of the 2019 Conference on Empirical Methods in
  Natural Language Processing and the 9th International Joint Conference on
  Natural Language Processing (EMNLP-IJCNLP)}, pages 3615--3620, Hong Kong,
  China. Association for Computational Linguistics.

\bibitem[{Brown et~al.(2020)Brown, Mann, Ryder, Subbiah, Kaplan, Dhariwal,
  Neelakantan, Shyam, Sastry, Askell, Agarwal, Herbert-Voss, Krueger, Henighan,
  Child, Ramesh, Ziegler, Wu, Winter, Hesse, Chen, Sigler, Litwin, Gray, Chess,
  Clark, Berner, McCandlish, Radford, Sutskever, and Amodei}]{gpt3}
Tom~B. Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared Kaplan,
  Prafulla Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda
  Askell, Sandhini Agarwal, Ariel Herbert-Voss, Gretchen Krueger, Tom Henighan,
  Rewon Child, Aditya Ramesh, Daniel~M. Ziegler, Jeffrey Wu, Clemens Winter,
  Christopher Hesse, Mark Chen, Eric Sigler, Mateusz Litwin, Scott Gray,
  Benjamin Chess, Jack Clark, Christopher Berner, Sam McCandlish, Alec Radford,
  Ilya Sutskever, and Dario Amodei. 2020.
\newblock \href {http://arxiv.org/abs/2005.14165} {Language models are few-shot
  learners}.

\bibitem[{Clark et~al.(2020)Clark, Luong, Le, and Manning}]{electra}
Kevin Clark, Minh-Thang Luong, Quoc~V. Le, and Christopher~D. Manning. 2020.
\newblock \href {https://openreview.net/forum?id=r1xMH1BtvB} {Electra:
  Pre-training text encoders as discriminators rather than generators}.
\newblock In \emph{International Conference on Learning Representations}.

\bibitem[{Conneau et~al.(2020)Conneau, Khandelwal, Goyal, Chaudhary, Wenzek,
  Guzm{\'{a}}n, Grave, Ott, Zettlemoyer, and Stoyanov}]{xlm-r}
Alexis Conneau, Kartikay Khandelwal, Naman Goyal, Vishrav Chaudhary, Guillaume
  Wenzek, Francisco Guzm{\'{a}}n, Edouard Grave, Myle Ott, Luke Zettlemoyer,
  and Veselin Stoyanov. 2020.
\newblock \href {https://www.aclweb.org/anthology/2020.acl-main.747/}
  {Unsupervised cross-lingual representation learning at scale}.
\newblock In \emph{Proceedings of the 58th Annual Meeting of the Association
  for Computational Linguistics, {ACL} 2020, Online, July 5-10, 2020}, pages
  8440--8451. Association for Computational Linguistics.

\bibitem[{Devlin et~al.(2019)Devlin, Chang, Lee, and Toutanova}]{bert}
Jacob Devlin, Ming-Wei Chang, Kenton Lee, and Kristina Toutanova. 2019.
\newblock \href {https://doi.org/10.18653/v1/N19-1423} {{BERT}: Pre-training of
  deep bidirectional transformers for language understanding}.
\newblock In \emph{Proceedings of the 2019 Conference of the North {A}merican
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies, Volume 1 (Long and Short Papers)}, pages 4171--4186,
  Minneapolis, Minnesota. Association for Computational Linguistics.

\bibitem[{d'Hoffschmidt et~al.(2020)d'Hoffschmidt, Belblidia, Brendlé,
  Heinrich, and Vidal}]{fquad}
Martin d'Hoffschmidt, Wacim Belblidia, Tom Brendlé, Quentin Heinrich, and
  Maxime Vidal. 2020.
\newblock \href {http://arxiv.org/abs/2002.06071} {Fquad: French question
  answering dataset}.

\bibitem[{Dodge et~al.(2020)Dodge, Ilharco, Schwartz, Farhadi, Hajishirzi, and
  Smith}]{seedfinetuning}
Jesse Dodge, Gabriel Ilharco, Roy Schwartz, Ali Farhadi, Hannaneh Hajishirzi,
  and Noah Smith. 2020.
\newblock \href {http://arxiv.org/abs/2002.06305} {Fine-tuning pretrained
  language models: Weight initializations, data orders, and early stopping}.

\bibitem[{Fan et~al.(2020{\natexlab{a}})Fan, Grave, and
  Joulin}]{dropoutpruning}
Angela Fan, Edouard Grave, and Armand Joulin. 2020{\natexlab{a}}.
\newblock \href {https://openreview.net/forum?id=SylO2yStDr} {Reducing
  transformer depth on demand with structured dropout}.
\newblock In \emph{8th International Conference on Learning Representations,
  {ICLR} 2020, Addis Ababa, Ethiopia, April 26-30, 2020}. OpenReview.net.

\bibitem[{Fan et~al.(2020{\natexlab{b}})Fan, Stock, Graham, Grave, Gribonval,
  Jegou, and Joulin}]{quantnoise}
Angela Fan, Pierre Stock, Benjamin Graham, Edouard Grave, Remi Gribonval, Herve
  Jegou, and Armand Joulin. 2020{\natexlab{b}}.
\newblock \href {http://arxiv.org/abs/2004.07320} {Training with quantization
  noise for extreme model compression}.

\bibitem[{Ganesh et~al.(2020)Ganesh, Chen, Lou, Khan, Yang, Chen, Winslett,
  Sajjad, and Nakov}]{survey}
Prakhar Ganesh, Yao Chen, Xin Lou, Mohammad~Ali Khan, Yin Yang, Deming Chen,
  Marianne Winslett, Hassan Sajjad, and Preslav Nakov. 2020.
\newblock \href {http://arxiv.org/abs/2002.11985} {Compressing large-scale
  transformer-based models: A case study on bert}.

\bibitem[{Hinton et~al.(2015)Hinton, Vinyals, and Dean}]{distillation}
Geoffrey Hinton, Oriol Vinyals, and Jeffrey Dean. 2015.
\newblock \href {http://arxiv.org/abs/1503.02531} {Distilling the knowledge in
  a neural network}.
\newblock In \emph{NIPS Deep Learning and Representation Learning Workshop}.

\bibitem[{Jiao et~al.(2019)Jiao, Yin, Shang, Jiang, Chen, Li, Wang, and
  Liu}]{tinybert}
Xiaoqi Jiao, Yichun Yin, Lifeng Shang, Xin Jiang, Xiao Chen, Linlin Li, Fang
  Wang, and Qun Liu. 2019.
\newblock \href {http://arxiv.org/abs/1909.10351} {Tinybert: Distilling bert
  for natural language understanding}.

\bibitem[{Kudo and Richardson(2018)}]{sentencepiece}
Taku Kudo and John Richardson. 2018.
\newblock \href
  {http://dblp.uni-trier.de/db/conf/emnlp/emnlp2018-d.html#KudoR18}
  {Sentencepiece: A simple and language independent subword tokenizer and
  detokenizer for neural text processing.}
\newblock In \emph{EMNLP (Demonstration)}, pages 66--71. Association for
  Computational Linguistics.

\bibitem[{Lample and Conneau(2019)}]{xlm}
Guillaume Lample and Alexis Conneau. 2019.
\newblock \href {http://arxiv.org/abs/1901.07291} {Cross-lingual language model
  pretraining}.

\bibitem[{Lee et~al.(2019)Lee, Yoon, Kim, Kim, Kim, So, and Kang}]{biobert}
Jinhyuk Lee, Wonjin Yoon, Sungdong Kim, Donghyeon Kim, Sunkyu Kim, Chan~Ho So,
  and Jaewoo Kang. 2019.
\newblock \href {https://doi.org/10.1093/bioinformatics/btz682} {{BioBERT: a
  pre-trained biomedical language representation model for biomedical text
  mining}}.
\newblock \emph{Bioinformatics}, 36(4):1234--1240.

\bibitem[{Li et~al.(2020)Li, Wallace, Shen, Lin, Keutzer, Klein, and
  Gonzalez}]{trainthencompress}
Zhuohan Li, Eric Wallace, Sheng Shen, Kevin Lin, Kurt Keutzer, Dan Klein, and
  Joseph~E. Gonzalez. 2020.
\newblock \href {http://arxiv.org/abs/2002.11794} {Train large, then compress:
  Rethinking model size for efficient training and inference of transformers}.

\bibitem[{Liu et~al.(2019)Liu, Ott, Goyal, Du, Joshi, Chen, Levy, Lewis,
  Zettlemoyer, and Stoyanov}]{roberta}
Yinhan Liu, Myle Ott, Naman Goyal, Jingfei Du, Mandar Joshi, Danqi Chen, Omer
  Levy, Mike Lewis, Luke Zettlemoyer, and Veselin Stoyanov. 2019.
\newblock \href {http://arxiv.org/abs/1907.11692} {Roberta: {A} robustly
  optimized {BERT} pretraining approach}.
\newblock \emph{CoRR}, abs/1907.11692.

\bibitem[{Martin et~al.(2020)Martin, Muller, Ortiz~Su{\'a}rez, Dupont, Romary,
  de~la Clergerie, Seddah, and Sagot}]{camembert}
Louis Martin, Benjamin Muller, Pedro~Javier Ortiz~Su{\'a}rez, Yoann Dupont,
  Laurent Romary, {\'E}ric de~la Clergerie, Djam{\'e} Seddah, and Beno{\^\i}t
  Sagot. 2020.
\newblock \href {https://doi.org/10.18653/v1/2020.acl-main.645} {{C}amem{BERT}:
  a tasty {F}rench language model}.
\newblock In \emph{Proceedings of the 58th Annual Meeting of the Association
  for Computational Linguistics}, pages 7203--7219, Online. Association for
  Computational Linguistics.

\bibitem[{McCarley et~al.(2019)McCarley, Chakravarti, and
  Sil}]{structuredpruning}
J.~S. McCarley, Rishav Chakravarti, and Avirup Sil. 2019.
\newblock \href {http://arxiv.org/abs/1910.06360} {Structured pruning of a
  bert-based question answering model}.

\bibitem[{{Ortiz Su{\'a}rez} et~al.(2019){Ortiz Su{\'a}rez}, Sagot, and
  Romary}]{oscar}
Pedro~Javier {Ortiz Su{\'a}rez}, Beno{\^i}t Sagot, and Laurent Romary. 2019.
\newblock \href {https://doi.org/10.14618/ids-pub-9021} {Asynchronous pipelines
  for processing huge corpora on medium to low resource infrastructures}.
\newblock Proceedings of the Workshop on Challenges in the Management of Large
  Corpora (CMLC-7) 2019. Cardiff, 22nd July 2019, pages 9 -- 16, Mannheim.
  Leibniz-Institut f{\"u}r Deutsche Sprache.

\bibitem[{Radford et~al.(2019)Radford, Wu, Child, Luan, Amodei, and
  Sutskever}]{gpt2}
Alec Radford, Jeffrey Wu, Rewon Child, David Luan, Dario Amodei, and Ilya
  Sutskever. 2019.
\newblock \href {https://openai.com/blog/better-language-models/} {Language
  models are unsupervised multitask learners}.
\newblock \emph{OpenAI Blog}, 1(8):9.

\bibitem[{Raffel et~al.(2020)Raffel, Shazeer, Roberts, Lee, Narang, Matena,
  Zhou, Li, and Liu}]{t5}
Colin Raffel, Noam Shazeer, Adam Roberts, Katherine Lee, Sharan Narang, Michael
  Matena, Yanqi Zhou, Wei Li, and Peter~J. Liu. 2020.
\newblock \href {http://jmlr.org/papers/v21/20-074.html} {Exploring the limits
  of transfer learning with a unified text-to-text transformer}.
\newblock \emph{Journal of Machine Learning Research}, 21(140):1--67.

\bibitem[{Rajpurkar et~al.(2016)Rajpurkar, Zhang, Lopyrev, and Liang}]{squad}
Pranav Rajpurkar, Jian Zhang, Konstantin Lopyrev, and Percy Liang. 2016.
\newblock \href {https://doi.org/10.18653/v1/D16-1264} {{SQ}u{AD}: 100,000+
  questions for machine comprehension of text}.
\newblock In \emph{Proceedings of the 2016 Conference on Empirical Methods in
  Natural Language Processing}, pages 2383--2392, Austin, Texas. Association
  for Computational Linguistics.

\bibitem[{Sanh et~al.(2020)Sanh, Wolf, and Rush}]{movementpruning}
Victor Sanh, Thomas Wolf, and Alexander~M. Rush. 2020.
\newblock \href {http://arxiv.org/abs/2005.07683} {Movement pruning: Adaptive
  sparsity by fine-tuning}.

\bibitem[{Sennrich et~al.(2016)Sennrich, Haddow, and Birch}]{bpe}
Rico Sennrich, Barry Haddow, and Alexandra Birch. 2016.
\newblock \href {https://doi.org/10.18653/v1/P16-1162} {Neural machine
  translation of rare words with subword units}.
\newblock In \emph{Proceedings of the 54th Annual Meeting of the Association
  for Computational Linguistics (Volume 1: Long Papers)}, pages 1715--1725,
  Berlin, Germany. Association for Computational Linguistics.

\bibitem[{Shen et~al.(2019)Shen, Dong, Ye, Ma, Yao, Gholami, Mahoney, and
  Keutzer}]{qbert}
Sheng Shen, Zhen Dong, Jiayu Ye, Linjian Ma, Zhewei Yao, Amir Gholami,
  Michael~W. Mahoney, and Kurt Keutzer. 2019.
\newblock \href
  {http://dblp.uni-trier.de/db/journals/corr/corr1909.html#abs-1909-05840}
  {Q-bert: Hessian based ultra low precision quantization of bert.}
\newblock \emph{CoRR}, abs/1909.05840.

\bibitem[{Sun et~al.(2020)Sun, Yu, Song, Liu, Yang, and Zhou}]{mobilebert}
Zhiqing Sun, Hongkun Yu, Xiaodan Song, Renjie Liu, Yiming Yang, and Denny Zhou.
  2020.
\newblock \href {https://doi.org/10.18653/v1/2020.acl-main.195}
  {{M}obile{BERT}: a compact task-agnostic {BERT} for resource-limited
  devices}.
\newblock In \emph{Proceedings of the 58th Annual Meeting of the Association
  for Computational Linguistics}, pages 2158--2170, Online. Association for
  Computational Linguistics.

\bibitem[{Turc et~al.(2019)Turc, Chang, Lee, and Toutanova}]{wellread}
Iulia Turc, Ming-Wei Chang, Kenton Lee, and Kristina Toutanova. 2019.
\newblock \href
  {http://dblp.uni-trier.de/db/journals/corr/corr1908.html#abs-1908-08962}
  {Well-read students learn better: The impact of student initialization on
  knowledge distillation.}
\newblock \emph{CoRR}, abs/1908.08962.

\bibitem[{Vaswani et~al.(2017)Vaswani, Shazeer, Parmar, Uszkoreit, Jones,
  Gomez, Kaiser, and Polosukhin}]{attention}
Ashish Vaswani, Noam Shazeer, Niki Parmar, Jakob Uszkoreit, Llion Jones,
  Aidan~N Gomez, \L~ukasz Kaiser, and Illia Polosukhin. 2017.
\newblock \href
  {http://papers.nips.cc/paper/7181-attention-is-all-you-need.pdf} {Attention
  is all you need}.
\newblock In I.~Guyon, U.~V. Luxburg, S.~Bengio, H.~Wallach, R.~Fergus,
  S.~Vishwanathan, and R.~Garnett, editors, \emph{Advances in Neural
  Information Processing Systems 30}, pages 5998--6008. Curran Associates, Inc.

\bibitem[{Wang et~al.(2018)Wang, Singh, Michael, Hill, Levy, and Bowman}]{glue}
Alex Wang, Amanpreet Singh, Julian Michael, Felix Hill, Omer Levy, and Samuel
  Bowman. 2018.
\newblock \href {https://doi.org/10.18653/v1/W18-5446} {{GLUE}: A multi-task
  benchmark and analysis platform for natural language understanding}.
\newblock In \emph{Proceedings of the 2018 {EMNLP} Workshop {B}lackbox{NLP}:
  Analyzing and Interpreting Neural Networks for {NLP}}, pages 353--355,
  Brussels, Belgium. Association for Computational Linguistics.

\bibitem[{Wolf et~al.(2019)Wolf, Debut, Sanh, Chaumond, Delangue, Moi, Cistac,
  Rault, Louf, Funtowicz, and Brew}]{huggingface}
Thomas Wolf, Lysandre Debut, Victor Sanh, Julien Chaumond, Clement Delangue,
  Anthony Moi, Pierric Cistac, Tim Rault, Rémi Louf, Morgan Funtowicz, and
  Jamie Brew. 2019.
\newblock \href {http://arxiv.org/abs/1910.03771} {Huggingface's transformers:
  State-of-the-art natural language processing}.

\end{thebibliography}

\pdfoutput=1

\documentclass[11pt,a4paper]{article}
\usepackage[hyperref]{emnlp2020}
\usepackage{times}
\usepackage{latexsym}

\usepackage{subcaption}

\usepackage{microtype}

\aclfinalcopy 
\def\aclpaperid{1351} 

\title{On the importance of pre-training data volume for compact language models}

\author{Vincent Micheli \\
  Illuin Technology / EPFL \\
  \texttt{vincent.micheli@illuin.tech}
  \And
  Martin d'Hoffschmidt \\
  Illuin Technology \\
  \texttt{martin.dhoffschmidt@illuin.tech}
  \AND
  François Fleuret \\
  UNIGE \\
  \texttt{francois.fleuret@unige.ch}
  }

\date{}

\begin{document}
\maketitle
\begin{abstract}
Recent advances in language modeling have led to computationally intensive and resource-demanding state-of-the-art models. In an effort towards sustainable practices, we study the impact of pre-training data volume on compact language models. Multiple BERT-based models are trained on gradually increasing amounts of French text. Through fine-tuning on the French Question Answering Dataset (FQuAD), we observe that well-performing models are obtained with as little as 100 MB of text. In addition, we show that past critically low amounts of pre-training data, an intermediate pre-training step on the task-specific corpus does not yield substantial improvements.
\end{abstract}

\section{Introduction}

Over the past year, pre-trained language models have become the norm in Natural Language Processing. These large-scale Transformer-based \citep{attention} networks considerably advanced the state-of-the-art in language understanding \citep{bert} via a two-step process: self-supervised learning on a vast text corpus followed by fine-tuning on a specific downstream task. 

Following these advances, the ongoing trend has been to build bigger models with an ever-increasing amount of data \citep{roberta, t5, gpt2, gpt3}. However, pre-training models with billions of parameters over hundreds of gigabytes of text requires tremendous computational resources that only a few companies and institutions can afford. Besides, many languages and specific corpora (e.g. legal, scientific) are currently under-resourced. Hence, our goal is to explore model architectures and data volumes lowering the entry barrier to new research and practical applications. 

We conduct experiments on French corpora in order to release the first French compact language models and to illustrate the training process in another language than English. Furthermore, we consider the question answering task since compact models may find their purpose in low-latency/fault-tolerant information retrieval systems.

\section{Problem statement}
We intend to study the impact of pre-training data volume when training compact bidirectional Transformers \citep{bert}. We assume a scarce resources setting, both in terms of data and computing power. Two key aspects are explored:

\begin{itemize}
    \item The amount of pre-training data required to train high-performing compact language models.
    \item The importance of corpus-specific MLM before fine-tuning.
\end{itemize}

We use the French part of the OSCAR corpora \citep{oscar} for pre-training and the FQuAD dataset \footnote{\url{https://illuin-tech.github.io/FQuAD-explorer/}} \citep{fquad} for machine reading comprehension fine-tuning. Moreover, the models under consideration are based on the CamemBERT \citep{camembert} language model.

\section{Related work}

A wealth of work has recently been released \citep{survey} on compressing Transformer-based models \citep{attention, bert} through the pre-training of compact models \citep{wellread}, distillation \citep{distillation, tinybert, mobilebert}, pruning \citep{trainthencompress, structuredpruning, movementpruning, dropoutpruning} and quantization \citep{qbert, quantnoise}. Nevertheless, absolute performance is not the end goal of this study. Rather, we investigate the training process of compact models in the absence of larger ones to distillate or prune. Furthermore, \citet{movementpruning} acknowledge the difficulty of speeding up sparse models due to the absence of specialized hardware. Therefore, from an inference speed standpoint, it is currently preferable to train compact models.

Language models have been successfully pre-trained on domain-specific corpora \citep{scibert, biobert} and outperform their general-purpose counterparts on targetted downstream tasks. Still, training these models involved large datasets and computational resources out of reach for most.

Multilingual models \citep{bert, xlm, xlm-r} have been released to alleviate the need for language-specific pre-training. While they offer competitive results, they usually lag behind monolingual models and require larger architectures.

\citet{camembert} observed that large models did not improve on evaluation tasks when increasing the amount of pre-training data from 4 GB to 138 GB. They left as future work to question the need for large scale pre-training corpora with other model architectures and tasks.

\section{Datasets}

\subsection{OSCAR}

OSCAR \footnote{\url{https://oscar-corpus.com/}} \citep{oscar} is a large-scale multilingual open source collection of corpora obtained by language classification and filtering of the Common Crawl corpus \footnote{ \url{https://commoncrawl.org/about/}}. The whole French part amounts to 138 GB of text and it has already been used to train French language models \citep{camembert}. In this work, we only extract a sample of 4 GB of shuffled lines.

\subsection{FQuAD}

FQuAD \citep{fquad} is a recently introduced open source French native reading comprehension dataset. It consists of 60,000 questions and answers gathered on a set of 1,769 high-quality Wikipedia articles. In many aspects, it is the French equivalent of SQuAD 1.1 \citep{squad}. Given a question and a paragraph, the task consists in extracting from the paragraph the span of text answering the question.

We chose FQuAD as the fine-tuning dataset because it allows one to draw a direct parallel with its English counterpart \citep{fquad} and is one of the largest annotated French datasets. However, question answering is a notoriously difficult task for compact models \citep{structuredpruning}. While distillation has shown to improve their results on the GLUE benchmark \citep{glue} substantially, machine reading comprehension remains difficult to speed-up without incurring a significant drop in accuracy. 

\section{CamemBERT\textsubscript{SMALL}} \label{camembert-small}

\begin{table}
\centering
\begin{tabular}{lrl}
\hline \textbf{Model} & \textbf{Size} & \textbf{Time} \\ \hline
CamemBERT\textsubscript{SMALL} & 72 MB & 157 ms \\
CamemBERT\textsubscript{BASE} & 440 MB & 705 ms \\
CamemBERT\textsubscript{LARGE} & 1340 MB & 2376 ms \\
\hline
\end{tabular}
\caption{\label{inferencetime-table} Model size and inference time on an Intel Xeon 2.30GHz Quad core CPU with batch size 1 and max sequence length 384 tokens (average over 1000 samples).}
\end{table}

CamemBERT \citep{camembert} is a multi-layer bidirectional Transformer \citep{attention} with two architectures: base (12 layers, 768 hidden dimensions, 12 attention heads, 110M parameters) and large (24 layers, 1024 hidden dimensions, 16 attention heads, 355M parameters). It is very similar to RoBERTa \citep{roberta}. The main differences are the use of whole-word masking and SentencePiece tokenization \citep{sentencepiece} instead of subword-masking and byte-level Byte-Pair encoding \citep{bpe, gpt2}. RoBERTa itself improves upon BERT by aggregating several modifications on top of the original architecture such as removing the next sentence prediction task, dynamic masking, and training with larger batches on more data. 

We introduce CamemBERT\textsubscript{SMALL} \footnote{The pre-trained models are made available in the Hugging Face collection: \url{https://huggingface.co/illuin/lepetit}.}, a CamemBERT-based language model with a small architecture (12 layers, 256 hidden dimensions, 4 attention heads, 17M parameters). The main difference with the original CamemBERT lies in the use of subword-masking. Indeed, the authors later found out that whole-word masking had at best a marginal impact on downstream task performance.

Apart from inference speed and size considerations, two main factors explain this architectural choice:

\begin{itemize}
    \item This is the same architecture as ELECTRA\textsubscript{SMALL++} \citep{electra}, a recently released compact language model. Even though ELECTRA and CamemBERT differ in many regards (ELECTRA being trained as a discriminator rather than a generator), prior experiments conducted by \citet{electra} give us an acceptable set of hyperparameters when pre-training and fine-tuning the model.
    \item \citet{wellread}'s empirical results suggest that depth should be prioritized over width when pre-training compact models. 
\end{itemize} 

Table \ref{inferencetime-table} shows that CamemBERT\textsubscript{SMALL} is much smaller and faster than its larger siblings. In a plausible setup for question answering systems, it provides, respectively, a 4.5-fold and 15-fold inference speed-up compared to CamemBERT\textsubscript{BASE} and CamemBERT\textsubscript{LARGE} while being 6.2 and 18.8 times smaller.

\section{Experiments} \label{experiments}

\begin{table}
\centering
\begin{tabular}{lll}
\hline \textbf{Hyperparameter} & \textbf{Pre-train} & \textbf{Fine-tune} \\ \hline
Train steps & 200k & 30k \\
Warmup steps & 10k & 3k \\
Batch size & 128 & 32 \\
Learning rate & 1e-4 & 1e-4 \\
Adam $\beta_1$ & 0.9 & 0.9 \\
Adam $\beta_2$ & 0.999 & 0.999 \\
Weight decay & 0.01 & 0.0 \\
Max gradient norm & 1.0 & 1.0 \\
Dropout & 0.1 & 0.1 \\
Mask percent & 15 & n/a \\
Max sequence length & 512 & 384 \\
\hline
\end{tabular}
\caption{\label{hyperparameters-table} Pre-training and fine-tuning hyperparameters. In the corpus-specific MLM step, we take the same hyperparameters as in pre-training except that we decrease the number of steps to 2.5k and drop the warmup.}
\end{table}

Six overlapping subsets are built from the 4 GB OSCAR sample. They are denoted as OSC\textsubscript{10}, OSC\textsubscript{100}, OSC\textsubscript{500}, OSC\textsubscript{1000}, OSC\textsubscript{2000} and OSC\textsubscript{4000} (the numbers indicating the number of MB). We extract an additional 10 MB sample from the corpus, which serves as a validation set for the self-supervised pre-training task. On the other hand, FQuAD consists of a train/dev split of 50,741 and 5,668 question/context pairs.

For each OSCAR subset, we pre-train a CamemBERT\textsubscript{SMALL} model with the standard masked language modeling (MLM) objective. 
Then we fine-tune the pre-trained models on the question answering task with the same span prediction method as BERT \citep{bert}. Between those two steps, an optional MLM step over the FQuAD train set is included.

Table \ref{hyperparameters-table} shows the pre-training, intermediate MLM (if any) and fine-tuning hyperparameters. 
Fine-tuning being a brittle process \citep{seedfinetuning}, fine-tuning results are averaged over 3 seeds. 

The experiments described were implemented using Hugging Face's Transformers library \citep{huggingface} and were conducted on an NVidia V100 16 GB.

\section{Analysis}

\citet{camembert} observed that complex downstream tasks may require more pre-training steps. Since for each OSCAR subset the validation loss is still slowly decreasing after 200k steps, we assume that training longer might increase performance on the difficult question answering task. On the other hand, corpus-specific MLM fine-tuning quickly converged for all models. Table \ref{results-table} reports the entirety of the results.

\begin{table*}
\begin{subtable}{.5\textwidth}
\centering
\begin{tabular}{lll}
\hline \textbf{Subset}  & \textbf{Perplexities} & \textbf{F1 score} \\ \hline
OSC\textsubscript{10}  & 45.20 / 43.34 & 58.18 (0.60) \\
OSC\textsubscript{100}  & 14.22 / 11.91 & 68.50  (0.25) \\
OSC\textsubscript{500}  & 12.75 / 10.58 & 69.50  (0.41)\\
OSC\textsubscript{1000}  & 12.56 / 10.57 & 69.35  (0.64)\\
OSC\textsubscript{2000}  & 12.45 / 10.41 & 70.96  (0.66)\\
OSC\textsubscript{4000}  & 12.49 / 10.35 & 69.76  (0.61)\\
\hline
\end{tabular}
\caption{\label{results-table-without} Without MLM fine-tuning.}
\end{subtable}
\hfill
\begin{subtable}{.5\textwidth}
\centering
\begin{tabular}{lll}
\hline \textbf{Subset}  & \textbf{Perplexities} & \textbf{F1 score} \\ \hline
OSC\textsubscript{10}  & 40.31 / 18.95 & 62.33  (0.58) \\
OSC\textsubscript{100}  & 16.35 / 9.41 & 69.04  (0.16) \\
OSC\textsubscript{500}  & 15.09 / 8.77 & 70.25  (0.43) \\
OSC\textsubscript{1000}  & 14.74 / 8.83  & 69.84  (0.27) \\
OSC\textsubscript{2000}  & 14.72 / 8.75 & 70.71  (0.08) \\
OSC\textsubscript{4000}  & 14.90 / 8.68 & 69.84  (0.79) \\
\hline
\end{tabular}
\caption{\label{results-table-with} With MLM fine-tuning.}
\end{subtable}
\caption{\label{results-table} Dev OSCAR / FQuAD perplexities and FQuAD F1 score (average token overlap between predicted and ground truth answers) for each pre-training subset.}
\end{table*}

\subsection{How much data does one need to pre-train a compact language model?}

As we increase the amount of pre-training data, perplexity on the OSCAR dev set decreases in every instance but one (OSC\textsubscript{4000}). Nevertheless, aside from OSC\textsubscript{10}, discrepancies are small and the models show almost identical learning curves. OSC\textsubscript{10} is underperforming in terms of MLM perplexity and question answering F1 score when compared to larger subsets. However, past this smallest dataset, pre-training data volume does not exhibit any strong monotonic relationship with downstream performance. The only OSCAR subset displaying a noticeable performance gap is OSC\textsubscript{2000}, with a +2.46 average F1 score increase over OSCAR\textsubscript{100}. For anchoring, a randomly initialized CamemBERT\textsubscript{SMALL} model "fine-tuned" directly on the FQuAD train set achieves an F1 score of only 17.76, i.e. 40 F1 points less than OSC\textsubscript{10}. This result indicates that even if a small amount of pre-training data is available, one should not neglect that step.
Regarding larger architectures, CamemBERT\textsubscript{BASE} and CamemBERT\textsubscript{LARGE} models from \citet{camembert} obtain an F1 score of 88 and 92, respectively, after fine-tuning.

Due to computational constraints, we could not investigate smaller or larger datasets as well as a prolonged pre-training phase. It could be the case that for a 200k pre-training steps budget, data volume is not the bottleneck. In fact, additional training steps may be even more beneficial for larger datasets. Nonetheless, a preliminary experiment pushing the pre-training phase of CamemBERT\textsubscript{SMALL} on OSC\textsubscript{2000} to 300k steps revealed that while the MLM loss decreased, the F1 score on the downstream task did not improve.

\subsection{Is corpus-specific MLM beneficial?}

Again, we observe a contrast between OSC\textsubscript{10} and larger subsets. OSC\textsubscript{10} is the only pre-training dataset significantly improving on the downstream task (+4.15 F1) and experiencing a decrease in perplexity on both pre-training and fine-tuning data when complemented with an intermediate MLM step. However, this corpus-specific MLM step is not truly intermediate since FQuAD contexts contain 10MB of raw text. This implies a 2-fold increase in pre-training data rather than a specific domain adaptation step. Therefore, we turn our focus to larger subsets for the rest of this analysis.

In these cases, MLM fine-tuning results in a net FQuAD perplexity decrease at the cost of an OSCAR perplexity increase. Domain shift may be the root cause of this trade-off. Indeed, as there exists a mismatch between pre-training and fine-tuning sets, the language model has to adapt to the specificity of descriptive paragraphs. In addition, perplexity is higher on the OSCAR dev set than on the FQuAD one. This is most likely due to the difficulty of predicting masked words in an heterogeneous web-crawled dataset compared to a set of high quality Wikipedia articles. 

For every pre-training subset but one (OSC\textsubscript{2000}), MLM fine-tuning induced a slight F1 score increase on the downstream task. However, these gains are marginal with at most a +0.75 average F1 score increase in the case of OSC\textsubscript{500}. Additional experiments are required to consolidate these findings, especially on larger task-specific datasets such as scientific or legal corpora. In those instances, a greater domain shift would probably legitimate an intermediate MLM fine-tuning step.

\section{Conclusion}

We investigated the importance of pre-training data volume when training compact Transformer-based models. We made the observation that 100 MB
of raw text are sufficient to reach similar performance as with larger datasets on a question answering task, and that corpus-specific self-supervised learning does not bring significant improvements on that particular problem. These preliminary results pave the way for further experiments with other language models, various architectures and new downstream tasks.

\section*{Acknowledgments}

We gratefully thank Quentin Heinrich for his reviewing and helpful discussions. We also thank Illuin Technology for its technical support and funding.

\newpage

\bibliographystyle{acl_natbib}
\bibliography{paper}

\end{document}